# Surrogate to Poincaré inequalities on manifolds for dimension reduction in nonlinear feature spaces

### v1

@@ 1 Introduction
where $\mathcal{J}(g)=\mathbb{E}\left[\|\Pi^{\perp}_{\text{span}\{G\}}\nabla u(\mathbf{X})\|_{2}^{2}\right]$ and $C(\mathbf{X}|\mathcal{G}_{m})$ is a constant depending only on the law of $\mathbf{X}$ and the model class $\mathcal{G}_{m}$. The constant $C(\mathbf{X}|\mathcal{G}_{m})$ is defined via probabilistic Poincaré constants associated to the conditional random variables $\mathbf{X}|G^{T}\mathbf{X}=\mathbf{z}$ for $\mathbf{z}\in\mathbb{R}^{m}$. The fact that the support of the latter lies in the $(d-m)$-dimensional affine space $\mathbf{z}+\text{span}\{(\}G)^{\perp}$ allows to derive bounds on $C(\mathbf{X}|\mathcal{G}_{m})$ for several classical probability measures [23]. The other main advantage of this method is that, while minimizing $\mathcal{E}$ on $\mathbb{R}^{d\times m}$ is a challenging problem, minimizing $\mathcal{J}$ on $\mathbb{R}^{d\times m}$ is rather straightforward. Indeed, it is a quadratic functional, whose minimum is the sum of the $d-m$ smallest eigenvalues of the matrix $\mathbb{E}\left[\nabla u(\mathbf{X})\nabla u(\mathbf{X})^{T}\right]\in\mathbb{R}^{d\times d}$, and the minimizer is a matrix whose columns are the eigenvectors associated with the $m$ largest eigenvalues of that matrix.
@@
Although linear dimension reduction has several theoretical and practical advantages, the main limitation is that it may not efficiently detect important nonlinear features, as it is easy to see by considering $u()\mathbf{x})=\mapsto h(\|\mathbf{x}\|^{2}_{2})$ for some $h\in\mathcal{C}^{1}$. In this regard, there are insights for developing nonlinear dimension reduction methods, where the model class $\mathcal{G}_{m}$ contains nonlinear functions. Until recently, there were mainly nonlinear sufficient dimension reduction methods developed, see for example [26, 14, 15, 16], extending the linear sufficient dimension reduction methods (SIR, SAVE, DR). More recently, various works introduced gradient-based nonlinear dimension reduction methods, either in a heuristic manner [28, 7, 22], or using Poincaré inequalities [3, 25]. In particular in [3] it was shown that (1.2) still holds, under important assumptions, with
@@
$\mathcal{J}(g)=\mathbb{E}\left[\|\Pi^{\perp}_{\text{span}\{\nabla\}g(\mathbf{X})}\nabla u(\mathbf{X})\|^{2}_{2}\right]=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|^{2}_{2}\right]-\mathbb{E}\left[\|\Pi_{\text{span}\{\nabla\}g(\mathbf{X})}\nabla u(\mathbf{X})\|^{2}_{2}\right].$ (1.3)
@@
The constant $C(\mathbf{X}|\mathcal{G}_{m})$ is here defined via probabilistic Poincaré inequalities on manifolds, associated to the conditional random variables $\mathbf{X}|g(\mathbf{X})=\mathbf{z}$ for $\mathbf{z}\in\mathbb{R}^{m}$. Indeed, the support of the latter lies in $g^{-1}(\{\mathbf{z}\})$, which is no longer an affine space but rather a $(d-m)$-dimensional smooth submanifold of $\mathbb{R}^{d}$, under some assumptions on $g$. Note that $\Pi^{\perp}_{\text{span}\{\nabla\}g(\mathbf{X})}\nabla u(\mathbf{X})$ is actually the Riemanian gradient of $u_{|g^{-1}(\{g(\mathbf{X})\})}$ at point $\mathbf{X}$, associated to the euclidean metric in $\mathbb{R}^{d}$.
@@
This nonlinear setting comes with two main issues. The first is that the theoretical foundation of (1.2) is much weaker than for linear features because controlling, or simply ensuring finiteness, for $C(\mathbf{X}|\mathcal{G}_{m})$ is a truly challenging task. A first step in overcoming this problem has been proposed in [25] by considering diffeomorphisms-based feature maps $g:\mathbf{x}\mapsto(\varphi_{1}(\mathbf{x}),\cdots,\varphi_{m}(\mathbf{x}))$, with $\varphi$ a $\mathcal{C}^{1}$ diffeomorphism on $\mathbb{R}^{d}$. The second issue is that optimizing $\mathcal{J}$ over a class $\mathcal{G}_{m}$ of nonlinear functions can be a challenging task, although first and second order optimization algorithms are available. This is mainly due to the nonlinearity of the mapping $g\mapsto\Pi^{\perp}_{\text{span}\{\nabla\}g(\mathbf{x})}$. This second issue is the motivation for the present work, in which we derive new upper bounds that yield to straightforward optimization procedures.
@@ 2.1 Error bounds based on Poincaré inequalities
Hence there is an incentive to build $g$ such that $\nabla u(\mathbf{x})$ and $\text{span}\{\nabla\}g(\mathbf{x})$ are as aligned as possible. The authors in [3] propose to minimize the following cost function,
@@
where $\Pi_{\nabla g(\mathbf{x})}:=\Pi_{\text{span}\{\nabla\}g(\mathbf{x})}\in\mathbb{R}^{d\times d}$ denotes the orthogonal projector from $\mathbb{R}^{d}$ onto $\text{span}\{\nabla\}g(\mathbf{x})$. The cost function $\mathcal{J}(g)$ naturally appears when using Poincaré inequalities and provides certified error bounds for $g\in\mathcal{G}_{m}$ [27]. When $\mathcal{G}_{m}$ contains nonlinear functions, these error bounds involve Poincaré inequalities on Riemannian manifolds, which we define in Definition 2.1 as in [3, 25]. Deriving these inequalities involve concepts from differential geometry, and although a deep understanding of the latter is not required to read the current section, we recommend [1, Chapter 3] for a comprehensive introduction to differential geometry and its applications.
@@
$T_{x}\mathcal{M}_{\mathbf{z}}=\text{ker}\nabla g(\mathbf{x})^{T}=(\mathrm{span}\nabla g(\mathbf{x}))^{\perp}$
@@
It is important to note that $\mathcal{J}$ satisfies some invariance properties. Indeed, for any diffeomorphism $\psi:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, by the chain rule and invertibility of $\nabla\psi(\mathbf{x})$, it holds that $\text{span}\{\nabla\}(\psi\circ g(\mathbf{x}))=\text{span}\{\nabla\}g(\mathbf{x})$ and thus
@@ 5 The case of multiple features
In this section, we propose an extension of the surrogate we introduced in (4.1) to the general case $m\geq 1$. The main problem when trying to do so is that for $m>1$, $\text{span}\{\nabla\}u(\mathbf{X})$ and $\text{span}\{\nabla\}g(\mathbf{X})$ are no longer of the same dimension, thus we can no longer expect to relate $\|\Pi^{\perp}_{\nabla g(\mathbf{X})}\nabla u(\mathbf{x})\|_{2}$ with $\|\Pi^{\perp}_{\nabla u(\mathbf{x})}\nabla g(\mathbf{X})\|_{F}$ in a similar way as in Lemma 4.2. For this reason, the extension we propose relies on learning one feature at a time, in a coordinate descent or greedy approach.
@@ 5.2 Minimizing the surrogate
The projectors $\Pi^{\perp}_{v_{g,j}(\mathbf{X})}$ and $\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}$ commute since $v_{g,j}(\mathbf{X})\in\text{span}\{(\}\nabla g_{-j}(\mathbf{X}))^{\perp}$. The definition of $\mathcal{L}_{m,j}$ in (5.2) then yields the result. ∎
@@ 6.1 Setting
Implementation.
The cross-validation procedures as well as the Kernel ridge regression rely on the library sklearn [20]. The optimization on Grassmann manifolds rely on on the library pymanopt [24]. The orthonormal polynomials feature maps rely on the python library tensap [19]. Open-source implementation of the approach proposed in this work, as well as tutorials, are available at https://github.com/alexandre-pasco/tensap/tree/poincare_learning.
@@ Appendix C Proofs for Section 5
Conversely, if $\mathcal{L}(g)=0$, then the second equation implies that almost surely $\|v\|_{2}^{2}=0$ or $\|\Pi^{\perp}_{v}w\|_{2}^{2}=0$. If $\|v\|_{2}^{2}=0$ then $\nabla u(\mathbf{X})\in\text{span}\{(\}W)$, or in other words the gradient of $u$ was already aligned with the gradient of the $m-1$ features $g_{-j}$, thus $\mathcal{J}(g)=0$. Otherwise if $\|\Pi^{\perp}_{v}w\|_{2}^{2}=0$, by Assumption 4.3 we have $w\neq 0$ almost surely, thus it yields $\|\Pi^{\perp}_{w}v\|_{2}^{2}=0$, or in other words
@@
Since $v_{g,j}(\mathbf{X})\in\text{span}\{(\}\nabla g_{-j}(\mathbf{X}))^{\perp}$, the projectors $\Pi^{\perp}_{v_{g,j}(\mathbf{X})}$ and $\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}$ commute. Hence $\|\Pi^{\perp}_{v_{g,j}(\mathbf{X})}w_{g,j}(\mathbf{X})\|_{2}^{2}=\nabla g_{j}(\mathbf{X})^{T}\Pi^{\perp}_{v_{g,j}(\mathbf{X})}\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}\nabla g_{j}(\mathbf{X})$. Also, $\nabla g_{j}(\mathbf{X})=\nabla\Phi(\mathbf{X})G_{j}$, thus

### v2

@@ 1 Introduction
where $\mathcal{J}(g)=\mathbb{E}\left[\|\Pi^{\perp}_{\text{span}\{G\}}\nabla u(\mathbf{X})\|_{2}^{2}\right]$ and $C(\mathbf{X}|\mathcal{G}_{m})$ is a constant depending only on the law of $\mathbf{X}$ and the model class $\mathcal{G}_{m}$. The constant $C(\mathbf{X}|\mathcal{G}_{m})$ is defined via probabilistic Poincaré constants associated to the conditional random variables $\mathbf{X}|G^{T}\mathbf{X}=\mathbf{z}$ for $\mathbf{z}\in\mathbb{R}^{m}$. The fact that the support of the latter lies in the $(d-m)$-dimensional affine space $\mathbf{z}+\text{span}\{G\}^{\perp}$ allows to derive bounds on $C(\mathbf{X}|\mathcal{G}_{m})$ for several classical probability measures [24]. The other main advantage of this method is that, while minimizing $\mathcal{E}$ on $\mathbb{R}^{d\times m}$ is a challenging problem, minimizing $\mathcal{J}$ on $\mathbb{R}^{d\times m}$ is rather straightforward. Indeed, it is a quadratic functional, whose minimum is the sum of the $d-m$ smallest eigenvalues of the matrix $\mathbb{E}\left[\nabla u(\mathbf{X})\nabla u(\mathbf{X})^{T}\right]\in\mathbb{R}^{d\times d}$, and the minimizer is a matrix whose columns are the eigenvectors associated with the $m$ largest eigenvalues of that matrix.
@@
Although linear dimension reduction has several theoretical and practical advantages, the main limitation is that it may not efficiently detect important nonlinear features, as it is easy to see by considering $u(\mathbf{x})=h(\|\mathbf{x}\|^{2}_{2})$ for some $h\in\mathcal{C}^{1}$. In this regard, there are insights for developing nonlinear dimension reduction methods, where the model class $\mathcal{G}_{m}$ contains nonlinear functions. Until recently, there were mainly nonlinear sufficient dimension reduction methods developed, see for example [27, 14, 15, 16], extending the linear sufficient dimension reduction methods (SIR, SAVE, DR). More recently, various works introduced gradient-based nonlinear dimension reduction methods, either in a heuristic manner [29, 7, 23], or using Poincaré inequalities [3, 26]. In particular in [3] it was shown that (1.2) still holds, under important assumptions, with
@@
$\mathcal{J}(g)=\mathbb{E}\left[\|\Pi^{\perp}_{\text{span}\{\nabla g(\mathbf{X})\}}\nabla u(\mathbf{X})\|^{2}_{2}\right]=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|^{2}_{2}\right]-\mathbb{E}\left[\|\Pi_{\text{span}\{\nabla g(\mathbf{X})\}}\nabla u(\mathbf{X})\|^{2}_{2}\right].$ (1.3)
@@
The constant $C(\mathbf{X}|\mathcal{G}_{m})$ is here defined via probabilistic Poincaré inequalities on manifolds, associated to the conditional random variables $\mathbf{X}|g(\mathbf{X})=\mathbf{z}$ for $\mathbf{z}\in\mathbb{R}^{m}$. Indeed, the support of the latter lies in $g^{-1}(\{\mathbf{z}\})$, which is no longer an affine space but rather a $(d-m)$-dimensional smooth submanifold of $\mathbb{R}^{d}$, under some assumptions on $g$. Note that $\Pi^{\perp}_{\text{span}\{\nabla g(\mathbf{X})\}}\nabla u(\mathbf{X})$ is actually the Riemanian gradient of $u_{|g^{-1}(\{g(\mathbf{X})\})}$ at point $\mathbf{X}$, associated to the euclidean metric in $\mathbb{R}^{d}$.
@@
This nonlinear setting comes with two main issues. The first is that the theoretical foundation of (1.2) is much weaker than for linear features because controlling, or simply ensuring finiteness, for $C(\mathbf{X}|\mathcal{G}_{m})$ is a truly challenging task. A first step in overcoming this problem has been proposed in [26] by considering diffeomorphisms-based feature maps $g:\mathbf{x}\mapsto(\varphi_{1}(\mathbf{x}),\cdots,\varphi_{m}(\mathbf{x}))$, with $\varphi$ a $\mathcal{C}^{1}$ diffeomorphism on $\mathbb{R}^{d}$. The second issue is that optimizing $\mathcal{J}$ over a class $\mathcal{G}_{m}$ of nonlinear functions can be a challenging task, although first and second order optimization algorithms are available. This is mainly due to the nonlinearity of the mapping $g\mapsto\Pi^{\perp}_{\text{span}\{\nabla g(\mathbf{x})\}}$. This second issue is the motivation for the present work, in which we derive new upper bounds that yield to straightforward optimization procedures.
@@ 2.1 Error bounds based on Poincaré inequalities
Hence there is an incentive to build $g$ such that $\nabla u(\mathbf{x})$ and $\text{span}\{\nabla{g(\mathbf{x})}\}$ are as aligned as possible. The authors in [3] propose to minimize the following cost function,
@@
where $\Pi_{\nabla g(\mathbf{x})}:=\Pi_{\text{span}\{\nabla g(\mathbf{x})\}}\in\mathbb{R}^{d\times d}$ denotes the orthogonal projector from $\mathbb{R}^{d}$ onto $\text{span}\{\nabla g(\mathbf{x})\}$. The cost function $\mathcal{J}(g)$ naturally appears when using Poincaré inequalities and provides certified error bounds for $g\in\mathcal{G}_{m}$ [28]. When $\mathcal{G}_{m}$ contains nonlinear functions, these error bounds involve Poincaré inequalities on Riemannian manifolds, which we define in Definition 2.1 as in [3, 26]. Deriving these inequalities involve concepts from differential geometry, and although a deep understanding of the latter is not required to read the current section, we recommend [1, Chapter 3] for a comprehensive introduction to differential geometry and its applications.
@@
$T_{\mathbf{x}}\mathcal{M}_{\mathbf{z}}=\text{ker}\{\nabla g(\mathbf{x})^{T}\}=\text{span}\{\nabla g(\mathbf{x})\}^{\perp}$
@@
It is important to note that $\mathcal{J}$ satisfies some invariance properties. Indeed, for any diffeomorphism $\psi:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, by the chain rule and invertibility of $\nabla\psi(\mathbf{x})$, it holds that $\text{span}\{\nabla(\psi\circ g(\mathbf{x}))\}=\text{span}\{\nabla g(\mathbf{x})\}$ and thus
@@ 5 The case of multiple features
In this section, we propose an extension of the surrogate we introduced in (4.1) to the general case $m\geq 1$. The main problem when trying to do so is that for $m>1$, $\text{span}\{\nabla u(\mathbf{X})\}$ and $\text{span}\{\nabla g(\mathbf{X})\}$ are no longer of the same dimension, thus we can no longer expect to relate $\|\Pi^{\perp}_{\nabla g(\mathbf{X})}\nabla u(\mathbf{x})\|_{2}$ with $\|\Pi^{\perp}_{\nabla u(\mathbf{x})}\nabla g(\mathbf{X})\|_{F}$ in a similar way as in Lemma 4.2. For this reason, the extension we propose relies on learning one feature at a time, in a coordinate descent or greedy approach.
@@ 5.2 Minimizing the surrogate
The projectors $\Pi^{\perp}_{v_{g,j}(\mathbf{X})}$ and $\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}$ commute since $v_{g,j}(\mathbf{X})\in\text{span}\{\nabla g_{-j}(\mathbf{X})\}^{\perp}$. The definition of $\mathcal{L}_{m,j}$ in (5.2) then yields the result. ∎
@@ 6.1 Setting
Implementation and data statement.
The cross-validation procedures as well as the Kernel ridge regression rely on the library sklearn [21]. The optimization on Grassmann manifolds rely on on the library pymanopt [25]. The orthonormal polynomials feature maps rely on the python library tensap [19]. The code underlying this article is freely available in [20] at https://doi.org/10.5281/zenodo.15430309.
@@ Appendix C Proofs for Section 5
Conversely, if $\mathcal{L}(g)=0$, then the second equation implies that almost surely $\|v\|_{2}^{2}=0$ or $\|\Pi^{\perp}_{v}w\|_{2}^{2}=0$. If $\|v\|_{2}^{2}=0$ then $\nabla u(\mathbf{X})\in\text{span}\{W\}$, or in other words the gradient of $u$ was already aligned with the gradient of the $m-1$ features $g_{-j}$, thus $\mathcal{J}(g)=0$. Otherwise if $\|\Pi^{\perp}_{v}w\|_{2}^{2}=0$, by Assumption 4.3 we have $w\neq 0$ almost surely, thus it yields $\|\Pi^{\perp}_{w}v\|_{2}^{2}=0$, or in other words
@@
Since $v_{g,j}(\mathbf{X})\in\text{span}\{\nabla g_{-j}(\mathbf{X})\}^{\perp}$, the projectors $\Pi^{\perp}_{v_{g,j}(\mathbf{X})}$ and $\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}$ commute. Hence $\|\Pi^{\perp}_{v_{g,j}(\mathbf{X})}w_{g,j}(\mathbf{X})\|_{2}^{2}=\nabla g_{j}(\mathbf{X})^{T}\Pi^{\perp}_{v_{g,j}(\mathbf{X})}\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}\nabla g_{j}(\mathbf{X})$. Also, $\nabla g_{j}(\mathbf{X})=\nabla\Phi(\mathbf{X})G_{j}$, thus